OpenStudy (anonymous):

a student must have the average (the mean) on five tests that is greater than or equal to 80% but less than 90% to receive a final grade of B. Devons grades on the first four tests were 70%, 65%, 90% and 77%. What range of grades on the fifth test would give him a B in the course? (Assume that 100% is the highest grade possible). Enter the range of grades necessary for Devon to earn a B average. _____% less than or greater to x less than or greater to 100%.

OpenStudy (anonymous):

Shouldn't it be 98%?

OpenStudy (anonymous):

oh yeah.. sorry..

OpenStudy (anonymous):

\[\Large \begin{array}{l} \frac{{70 + 65 + 90 + 77 + x}}{5} \ge 80\\ \frac{{302}}{5} + \frac{x}{5} \ge 80\\ 302 + x \ge 400\\ x \ge 98 \end{array}\]

OpenStudy (anonymous):

\[80\leq \frac{65+70+77+90+x}{5}<90\] is your equation

OpenStudy (anonymous):

probably my brain is not working correctly

OpenStudy (anonymous):

Heh mine's never working correctly, especially when tired.

OpenStudy (anonymous):

actually it is not an equation it is an inequality. more coffee!

OpenStudy (anonymous):

To the coffee pot!

OpenStudy (anonymous):

so is the answer 98%

OpenStudy (anonymous):

I need to enter the range of grades

OpenStudy (anonymous):

\[80\leq \frac{302+x}{5}<90\] \[400\leq 302+x<450\] \[98\leq x<142\] guy doesn't have a chance and should try for a C+

OpenStudy (anonymous):

\[x\geq 98\]

OpenStudy (anonymous):

btw i hope this is a typo less than or greater to x less than or greater to 100%. there is no such thing is "less than or greater to"

OpenStudy (anonymous):

i meant equals sorry

OpenStudy (anonymous):

98% is not the correct answer

OpenStudy (anonymous):

It has to be. His final score must be equal to or greater than 98% to get an average of 80% or more.